# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v2

@@ 3.2 Identification of Latent Variables
In this subsection, we demonstrate that given further sparsity assumptions, the latent Markov network over ${\mathbf{c}}_{t}$ and the latent variables are also identifiable. For a better explanation of the identifiability of latent variables, we first introduce the definition of the Intimate Neighbor Set [36].
Definition 2 (Intimate Neighbor Set [36]).
@@
$\small\begin{split}\Psi_{\mathcal{M}_{{\mathbf{c}}_{t}}}\!(c_{t,i})\!\triangleq\!\{c_{t,j}|c_{t,j}\text{ is adjacent to }c_{t,i},\text{ and it is also adjacent to all other neighbors of }c_{t,i},c_{t,j}\!\in\!{\mathbf{c}}_{t}\!\backslash\!\{c_{t,i}\}\!\}\end{split}$
@@
(Component-wise Identification of Latent Variables with instantaneous dependencies.) Suppose that the observations are generated by Equation (1) and (2), and $\mathcal{M}_{{\mathbf{c}}_{t}}$ is the Markov network over ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t},{\mathbf{z}}_{t+1}\}$. Except for the assumptions A1 and A2 from Theorem 1, we further make the following assumption:
  • A3 (Latent Process Sparsity): For any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is an empty set.
@@
Proof Sketch. Given the fact that there always exists a row permutation for each invertible matrix such that the permuted diagonal entries are non-zero [33], we utilize the conclusion of Theorem 1 to demonstrate that any edge present in the true Markov network will necessarily exist in the estimated one. Furthermore, when the number of estimated edges reaches a minimum, the identifiability of Markov network can be achieved. Finally, we once again leverage Theorem 1 to illustrate that the permutation which allows the Markov network to be identified can further lead to a component-wise level identifiability of latent variables. A detailed proof is given in Appendix A2.3.
Discussion. Intuitively, Theorem 2 tells us that any latent variable $z_{t,i}$ is identifiable when its intimate set is empty, which benefits from the sparse causal influence. Compared with the existing identifiability results for instantaneous dependencies like intervention or grouping observations, this assumption is more practical and reasonable in real-world time series datasets. For instance, in human motion forecasting, the human skeleton is a sparse structure, since the joints in arms and legs are significantly independent even if instantaneous dependencies exist in adjacent joints.
Figure 2: Examples of Markov networks with different types of contextual information that satisfy the sparse latent process assumption. Case (i) uses both historical and future information for identifiability. Case (ii) uses only historical information. Case (iii) that does not include instantaneous dependencies can identify latent variables without any contextual information, which degenerates to TDRL[39]. Please note that in all cases, we show only the Markov net of ${\mathbf{c}}_{t}$ conditioned on previous timestamps while omitting the conditions for simplicity.
@@
(General Case for Component-wise Identification.) Suppose that the observations are generated by Equation (1) and (2), and there exists ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-a},\cdots,{\mathbf{z}}_{t},\cdots,{\mathbf{z}}_{t+b}\}$ with the corresponding Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$. Suppose assumptions A1 and A2 hold true, and for any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is an empty set. When the observational equivalence is achieved with the minimal number of edges of estimated Markov network of $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$, there must exist a permutation $\pi$ of the estimated latent variables, such that $z_{t,i}$ and $\hat{z}_{t,\pi(i)}$ is one-to-one corresponding, i.e., $z_{t,i}$ is component-wise identifiable.
@@
Discussion. In this part, we give a further discussion on how contextual information can help to identify current latent variables. Let us take Figure 2 (i) as an example, when we consider the current latent variables $z_{t,1},z_{t,2}$ and $z_{t,3}$ only, it is easy to see that neither $z_{t,1}$ nor $z_{t,3}$ are identifiable since both of their intimate set are $\{z_{t,2}\}$. However, when the historical information ${\mathbf{z}}_{t-1}$ is taken into consideration, where $z_{t-1,1}$ affects only $z_{t,1}$ but not $z_{t,2}$, then $z_{t,1}$ will immediately be identified since the intimate set of $z_{t,1}$ will be empty thanks to the unique influence from $z_{t-1,1}$. Similarly, with a future variable $z_{t+1,3}$ which is influenced by $z_{t,3}$ but not $z_{t,1}$, then $z_{t,3}$ can be identified as well. Besides, under some circumstances, only part of the contextual information is needed. Therefore, the length of the required time window $L$ can be reduced as well, which leads to a more relaxed requirement of assumption A2, according to Theorem 1. As shown in Figure 2 (ii), with only historical information, the identifiability of ${\mathbf{z}}_{t}$ can already be achieved. Furthermore, some existing identifiability results [23, 39] can be considered as a special case of our theoretical results when the instantaneous effects vanish. For example, Figure 2 (iii) provides the example of TDRL [39], where all intimate sets are empty naturally, and the identifiability is assured.
@@ 4.1 Temporally Variational Inference Architecture
Figure 3: The framework of the IDOL model. The encoder and decoder are used for latent variables extraction and observation reconstruction. The prior network is used for prior distribution estimation, and $L_{s}$ denotes the gradient-based sparsity penalty. The solid and dashed arrows denote the forward and backward propagation, respectively.
Based on the data generation process, we derive the evidence lower bound (ELBO) as shown in Equation (6). Appendix A3.2 illustrates the derivation details of the ELBO.
$\small\begin{split}ELBO=&\underbrace{\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\ln p({\mathbf{x}}_{1:T|}|{\mathbf{z}}_{1:T})}_{L_{r}}\\ &-\alpha\underbrace{D_{KL}(q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})||p({\mathbf{z}}_{1:T}))}_{L_{K}},\end{split}$ (7)
where $\alpha$ denotes the hyper-parameters and $D_{KL}$ denotes the KL divergence. $q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})$ and $p({\mathbf{x}}_{1:T|}|{\mathbf{z}}_{1:T})$ which are used to approximate the prior distribution and reconstruct the observations. Formally, the encoder and the decoder can be formalized as follows.
$\small{\mathbf{z}}_{1:T}=\phi({\mathbf{x}}_{1:T}),\quad\hat{{\mathbf{x}}}=\psi({\mathbf{z}}_{1:T}),$ (8)
where $\phi$ and $\psi$ denote the neural architecture based on convolution neural networks. Based on theoretical results, we develop the IDOL model as shown in Figure 3, which is built on the variational inference to model the distribution of observations. To estimate the prior distribution and enforce the independent noise assumption, we devise the prior networks. Moreover, we employ a gradient-based sparsity penalty to promote the sparse causal influence.
@@ 4.2 Prior Estimation
To estimate the prior distribution of latent variables, we propose the prior networks. Specifically, we first let ${r_{i}}$ be a set of learned inverse transition functions that take the estimated latent variables as input to estimate the noise term $\hat{\epsilon}_{i}$, i.e. $\hat{\epsilon}_{t,i}=r_{i}(\hat{{\mathbf{z}}}_{t},\hat{{\mathbf{z}}}_{t-1})$ [^1] We use the superscript symbol to denote estimated latent variables. And each $r_{i}$ is implemented by Multi-layer Perceptron networks (MLPs). Sequentially, we devise a transformation $\kappa:=\{\hat{{\mathbf{z}}}_{t-1},\hat{{\mathbf{z}}}_{t}\}\rightarrow\{\hat{{\mathbf{z}}}_{t-1},\hat{\bm{\epsilon}}_{t}\}$, whose Jacobian can be formalized as ${\mathbf{J}_{\kappa}=\begin{pmatrix}\mathbb{I}&0\\ \mathbf{J}_{d}&\mathbf{J}_{e}\end{pmatrix}}$, where $\mathbf{J}_{d}\!\!=\!\!\text{diag}\!\left(\frac{\partial r_{i}}{\partial\hat{z}_{t-1,i}}\!\right)$ and $\mathbf{J}_{e}\!\!=\!\!\text{diag}\!\left(\frac{\partial r_{i}}{\partial\hat{z}_{t,i}}\!\right)$. Hence we have Equation (8) via the change of variables formula.
$\small\log p(\hat{{\mathbf{z}}}_{t},\hat{{\mathbf{z}}}_{t-1})=\log p(\hat{{\mathbf{z}}}_{t-1},\epsilon_{t})+\log|\frac{\partial r_{i}}{\partial z_{t,i}}|.$ (9)
@@
$\small\log p(\hat{{\mathbf{z}}}_{1:T})=p(\hat{z}_{1})\prod_{\tau=2}^{T}\left(\sum_{i=1}^{n}\log p(\hat{\epsilon}_{\tau,i})+\sum_{i=1}^{n}\log|\frac{\partial r_{i}}{\partial z_{t,i}}|\right),$ (10)
where $p(\hat{\epsilon}_{\tau,i})$ is assumed to follow a Gaussian distribution. Please refer to Appendix A3.1 for more details of the prior derivation.
@@ 4.3 Gradient-based Sparsity Regularization
Ideally, the MLPs-based architecture $r_{i}$ can capture the causal structure of latent variables by restricting the independence of noise $\hat{\epsilon}_{t,i}$. However, without any further constraint, $r_{i}$ may bring redundant causal edges from $\𝑖{\hat{{\mathbf{z}}}_{t-1},\hat{{\mathbf{z}}}_{t,[n]\textbackslash i}}$ to $\hat{z}_{t,i}$, leading to the incorrect estimation of prior distribution and further the suboptimization of ELBO. As mentioned in Subsection 4.2, $\mathbf{J}_{d}$ and $\mathbf{J}_{e}$ intuitively denote the time-delayed and instantaneous causal structures of latent variables, since they describe how the $\𝑖{\hat{{\mathbf{z}}}_{t-1},\hat{{\mathbf{z}}}_{t,[n]\textbackslash i}}$ contribute to $\hat{z}_{t,i}$, which motivate us to remove these redundant causal edges with a sparsity regularization term $\mathcal{L}_{S}$ by simply applying the $\mathcal{L}_{1}$ on $\mathbf{J}_{d}$ and $\mathbf{J}_{e}$. Formally, we have
$\small L_{S}=||\mathbf{J}_{d}||_{1}+||\mathbf{J}_{e}||_{1}.,$ (11)
@@
$\small\mathcal{L}_{total}=-\mathcal{L}_{r}-\alpha\mathcal{L}_{K}+\beta\mathcal{L}_{S},$ (12)
@@ 5.1.1 Experimental Setup
Data Generation. We generate the simulated time series data with the fixed latent causal process as introduced in Equations (1)-(2) and Figure 1 (c). To better evaluate the proposed theoretical results, we provide three datasets with an increasing sparsity degree, which are named Dataset A, B, and C, respectively. We further generate a dense dataset with a dense latent process. Please refer to Appendix A4.1.1 for the details of data generation and evaluation metrics.
Baselines. To evaluate the effectiveness of our method, we consider the following compared methods. First, we consider the standard $\beta$-VAE [40] and FactorVAE [41], which ignores historical information and auxiliary variables. Moreover, we consider TDRL [39], iVAE [24], TCL [18], PCL [30], and SlowVAE [42], which use temporal information but do not assume instantaneous dependency on latent processes. Finally, we consider the iCITRIS [34] and G-CaRL [35], which are devised for latent causal processes with instantaneous dependencies but require interventions or grouping of observations. As for G-CaRL, we follow the implementation description in the original paper and assign random grouping for observations. As for iCITRIS, we assign a random intervention variable for latent variables. We repeat each method over three random seeds and report the average results.

### v3

@@ 3.2 Identification of Latent Variables
In this subsection, we demonstrate that given further sparsity assumptions, the latent Markov network over ${\mathbf{c}}_{t}$ and the latent variables are also identifiable. For a better explanation of the identifiability of latent variables, we first introduce the definition of the Intimate Neighbor Set (zhang2024causal).
Definition 2 (Intimate Neighbor Set (zhang2024causal)).
@@
[…]
In other words, $\Psi$ contains all neighbors, iff all neighbors of $c_{t,i}$ are also in the same unique clique as $c_{t,i}$, else $\Psi=\emptyset$. We further provide an example in the Appendix B.8.
@@
(Component-wise Identification of Latent Variables with instantaneous dependencies.) Suppose that the observations are generated instantaneous latent process, and $\mathcal{M}_{{\mathbf{c}}_{t}}$ is the Markov network over ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t},{\mathbf{z}}_{t+1}\}\in\mathbb{R}^{3n}$. Except for the smooth, positive density and sufficient variability assumptions, we further make the following assumption:
  • A3 (Sparse Latent Process): 1) For any $z_{it}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{it}$ is an empty set. 2) For any $z_{it}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{it}$ is not an empty set, but there exists a direct child $z_{jt}$, whose intimate neighbor set is an empty set.
@@
(i) The estimated Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ is isomorphic  [^1] Please refer to the definition of isomorphism of Markov networks in Appendix B.7. to the ground-truth Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$.
(ii) There exists a permutation $\pi$ of the estimated latent variables, such that $z_{it}$ and $\hat{z}_{\pi(i)t}$ is one-to-one corresponding, i.e., $z_{it}$ is component-wise identifiable.
Figure 2: Examples of Markov networks with different types of contextual information that satisfy the sparse latent process assumption. Case (i) uses both historical and future information for identifiability. Case (ii) uses only historical information. Case (iii) that does not include instantaneous dependencies can identify latent variables without any contextual information, which degenerates to TDRL (yao2022temporally). Please note that in all cases, we show only the Markov net of ${\mathbf{c}}_{t}$ conditioned on previous timestamps while omitting the conditions for simplicity.
Proof Sketch. Given the fact that there always exists a row permutation for each invertible matrix such that the permuted diagonal entries are non-zero (zheng2024generalizing), we utilize the conclusion of Theorem 1 to demonstrate that any edge present in the true Markov network will necessarily exist in the estimated one. Furthermore, when the number of estimated edges reaches a minimum, the identifiability of Markov network can be achieved. Finally, we once again leverage Theorem 1 to illustrate that the permutation which allows the Markov network to be identified can further lead to a component-wise level identifiability of latent variables. The detailed proof and discussion of assumptions are given in Appendix B.3 and C, respectively.
Discussion. Intuitively, Theorem 2 tells us that any latent variable $z_{t,i}$ is identifiable when its intimate set is empty, which benefits from the sparse causal influence. Compared with the existing identifiability results for instantaneous dependencies like intervention or grouping observations, this assumption is more practical and reasonable in real-world time series datasets.
@@
(General Case for Component-wise Identification.) Suppose that the observations are generated by Equation (1) and (2), and there exists ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-a},\cdots,{\mathbf{z}}_{t},\cdots,{\mathbf{z}}_{t+b}\}\in\mathbb{R}^{(a+b+1)\times n}$ with the corresponding Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$. Suppose assumptions A1 and A2 hold true, and for any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is an empty set. When the observational equivalence is achieved with the minimal number of edges of the estimated Markov network of $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$, there must exist a permutation $\pi$ of the estimated latent variables, such that $z_{t,i}$ and $\hat{z}_{t,\pi(i)}$ is one-to-one corresponding, i.e., $z_{t,i}$ is component-wise identifiable.
@@
Discussion. In this part, we give a further discussion on how contextual information can help to identify current latent variables. Let us take Figure 2 (i) as an example, when we consider the current latent variables $z_{t,1},z_{t,2}$ and $z_{t,3}$ only, it is easy to see that neither $z_{t,1}$ nor $z_{t,3}$ are identifiable since both of their intimate set are $\{z_{t,2}\}$. However, when the historical information ${\mathbf{z}}_{t-1}$ is taken into consideration, where $z_{t-1,1}$ affects only $z_{t,1}$ but not $z_{t,2}$, then $z_{t,1}$ will immediately be identified since the intimate set of $z_{t,1}$ will be empty thanks to the unique influence from $z_{t-1,1}$. Similarly, with a future variable $z_{t+1,3}$ which is influenced by $z_{t,3}$ but not $z_{t,1}$, then $z_{t,3}$ can be identified as well. Besides, under some circumstances, only part of the contextual information is needed. Therefore, the length of the required time window $L$ can be reduced as well, which leads to a more relaxed requirement of assumption A2, according to Theorem 1. As shown in Figure 2 (ii), with only historical information, the identifiability of ${\mathbf{z}}_{t}$ can already be achieved. Furthermore, some existing identifiability results (yao2021learning; yao2022temporally) can be considered as a special case of our theoretical results when the instantaneous effects vanish. For example, Figure 2 (iii) provides the example of TDRL (yao2022temporally), where all intimate sets are empty naturally, and the identifiability is assured.
Finally, we explore under what circumstances the causal graph of latent variables is also identifiable.
Theorem 3.
(Identification of Latent Causal Process.) Suppose that the observations are generated by Equation (1)-(2), and that $\mathcal{M}_{{\mathbf{c}}_{t}}$ is the Markov network over ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t}\}\in\mathbb{R}^{2n}$. Suppose that all assumptions for Theorem 2 hold. We further make the following assumption: for any pair of adjacent latent variables $z_{t,i},z_{t,j}$ at time step $t$, their time-delayed parents are not identical, i.e., $\text{Pa}_{d}(z_{t,i})\not=\text{Pa}_{d}(z_{t,j})$. Then, the causal graph of the latent causal process is identifiable.
Proof Sketch. We first establish the identification of the skeleton of the causal graph using the results of Theorem 2. The directions of time-delayed edges are fully determined since time can only move forward. Finally, we leverage the v-structure to identify the instantaneous edges. A detailed proof is provided in Appendix B.5.
Table 1: The summary of related work of causal representation learning.
No Intervention No Grouping Stationarity Instantaneous Effect Temporal Data
IDOL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
yao2022temporally $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
morioka2023causal $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
lippe2023causal $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
zhang2024causal $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
3.3 Comparasion with Existing Methods
The application of real-world scenarios presents numerous new opportunities and challenges. Compared to existing methods of temporally causal representation learning with instantaneous effects, our approach significantly relaxes the constraints on theoretical assumptions, as shown in Table 1. TDRL (yao2022temporally) requires that there are no instantaneous effects during the transition process, which can be considered as a special case of our method. G-CaRL (morioka2023causal) and iCITRIS (lippe2023causal) require extra intervention or grouping information, which is not necessary in our method. zhang2024causal focus on non-temporal scenarios and require extra domain labels. Instead, IDOL can handle the stationary temporal case. Besides, when the instantaneous effect is dense, which is beyond the capacity of zhang2024causal, IDOL can still utilize the contextual information for better identifiability as long as the temporal transition is sparse.
@@ 4 Identifiable Instantaneous Latent Dynamic Model
Figure 3: The framework of the IDOL model. The encoder and decoder are used for the extraction of latent variables and observation reconstruction. The prior network is used for prior distribution estimation, and $L_{s}$ denotes the gradient-based sparsity penalty. The solid and dashed arrows denote the forward and backward propagation.
@@ 4.1 Temporally Variational Inference Architecture
Based on the data generation process, we derive the evidence lower bound (ELBO) as shown in Equation (7) for general time series modeling. Please refer to Appendix E.2 for the derivation details of the EBLO for time series forecasting.
$\color[rgb]{0,0,0}\begin{split}ELBO=&\underbrace{\mathbb{E}_{q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})}\ln p({\mathbf{x}}_{1:T|}|{\mathbf{z}}_{1:T})}_{\mathcal{L}_{r}}\\ &-\underbrace{D_{KL}(q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})||p({\mathbf{z}}_{1:T}))}_{\mathcal{L}_{K}},\end{split}$ (7)
where $D_{KL}$ denotes the KL divergence. We use posterior $q({\mathbf{z}}_{1:T}|{\mathbf{x}}_{1:T})$ to approximate the prior $p({\mathbf{z}}_{1:T})$. Besides, $p({\mathbf{x}}_{1:T}|{\mathbf{z}}_{1:T})$ is used to reconstruct the observations. The encoder and the decoder can be formalized as follows:
${\mathbf{z}}_{1:T}=\phi({\mathbf{x}}_{1:T}),\quad\hat{{\mathbf{x}}}=\psi({\mathbf{z}}_{1:T}),$ (8)
where $\phi$ and $\psi$ denote the neural architecture based on convolution neural networks. Please refer to Table A4 for the implementation details of $\phi$ and $\psi$.
Based on theoretical results, we develop the IDOL model as shown in Figure 3, which is built on the variational inference to model the distribution of observations. To estimate the prior distribution and enforce the independent noise assumption, we devise the prior networks. Moreover, we employ a gradient-based sparsity penalty to promote the sparse causal influence.
@@ 4.2 Prior Estimation
To enforce the independence of noise in Equation (2), we minimize the Kullback-Leibler (KL) divergence between the posterior distribution $\prod_{i}p(z_{t,i}|\text{Pa}_{d}(z_{t,i}),\text{Pa}_{e}(z_{t,i}),{\mathbf{x}}_{t})$ and a prior $\prod_{i}p(z_{t,i}|\text{Pa}_{d}(z_{t,i}),\text{Pa}_{e}(z_{t,i}))$. This minimization implies that latent variables are mutually independent and conditioned on their historical and instantaneous parents. However, since the prior distribution may have any arbitrary density function, it is difficult to estimate such a prior. To solve this challenge, we follow chen2024caring, yao2022temporally and propose the prior networks. Specifically, we first let $r_{i}$} be an inversed function of $f$ from Equation (2) that are implemented by normalizing flow. papamakarios2021normalizing take the estimated latent variables as input to estimate the noise term $\hat{\epsilon}_{i}$, i.e. $\hat{\epsilon}_{t,i}=r_{i}(\hat{{\mathbf{z}}}_{t},\hat{{\mathbf{z}}}_{t-1})$ [^2] We use the superscript symbol to denote estimated latent variables. And each $r_{i}$ is implemented by Multi-layer Perceptron networks (MLPs). Sequentially, we devise a transformation $\kappa:=\{\hat{{\mathbf{z}}}_{t-1},\hat{{\mathbf{z}}}_{t}\}\rightarrow\{\hat{{\mathbf{z}}}_{t-1},\hat{\bm{\epsilon}}_{t}\}$, whose Jacobian can be formalized as $\mathbf{J}_{\kappa}=\begin{pmatrix}\mathbb{I}&0\\ \mathbf{J}_{d}&\mathbf{J}_{e}\end{pmatrix}$}, where $\mathbf{J}_{d}=\left(\frac{\partial r_{i}}{\partial\hat{z}_{t-1,i}}\right)$ and $\mathbf{J}_{e}=\left(\frac{\partial r_{i}}{\partial\hat{z}_{t,i}}\right)$. Hence we have Equation (9) via the change of variables formula.
$\log p(\hat{{\mathbf{z}}}_{t},\hat{{\mathbf{z}}}_{t-1})=\log p(\hat{{\mathbf{z}}}_{t-1},\epsilon_{t})+\log|\frac{\partial r_{i}}{\partial z_{t,i}}|.$ (9)
@@
$\log p(\hat{{\mathbf{z}}}_{1:T})=\log p(\hat{z}_{1})+\sum_{\tau=2}^{T}\left(\sum_{i=1}^{n}\log p(\hat{\epsilon}_{\tau,i})+\sum_{i=1}^{n}\log|\frac{\partial r_{i}}{\partial z_{t,i}}|\right),$ (10)
where $p(\hat{\epsilon}_{\tau,i})$ is assumed to follow a Gaussian distribution. Please refer to Appendix E.1 for more details of the prior derivation.
@@ 4.3 Gradient-based Sparsity Regularization
[…]
@@
$\mathcal{L}_{total}=-\mathcal{L}_{r}-\alpha\mathcal{L}_{K}+\beta\mathcal{L}_{S},$ (12)
@@ 5.1.1 Experimental Setup
Data Generation. We generate the simulated time series data with the fixed latent causal process as introduced in Equations (1)-(2) and Figure 1 (c). To better evaluate the proposed theoretical results, we provide six synthetic datasets from A to F with 3, 5, 8, 8, 8, 16 latent variables, respectively. Dataset D contains no instantaneous effects, which degenerate to the TDRL (yao2022temporally) setting as introduced in Figure 2 (iii). Dataset E has a dense latent causal process that violates the assumption A3(Latent Process Sparsity). All datasets except E satisfy the assumption A3. Please refer to Appendix F.1.1 for the details of data generation and evaluation metrics.
Baselines. To evaluate the effectiveness of our method, we consider the following compared methods. First, we consider the standard $\beta$-VAE (higgins2017beta) and FactorVAE (kim2018disentangling), which ignores historical information and auxiliary variables. Moreover, we consider TDRL (yao2022temporally), iVAE (khemakhem2020variational), TCL (hyvarinen2016unsupervised), PCL (hyvarinen2017nonlinear), and SlowVAE (klindt2020towards), which use temporal information but do not assume instantaneous dependency on latent processes. Finally, we consider the iCITRIS (lippe2023causal) and G-CaRL (morioka2023causal), which are devised for latent causal processes with instantaneous dependencies but require interventions or grouping of observations. As for G-CaRL, we follow the implementation description in the original paper and assign random grouping for observations. As for iCITRIS, we assign a random intervention variable for latent variables. We repeat each method over three random seeds and report the average results.